zarkam21:

Susan is looking into a new cell phone carrier. She researches her options and finds the two companies which have the best reception in her neighborhood offer very different monthly plans. Write a function rule which represents each company’s monthly fees. A) Company A offers unlimited minutes and texts for a flat fee of $50 per month and charges $1.50 for each GB of data used per month. B) Company B offers unlimited minutes and texts for a flat fee of $35 per month and charges $5.00 for each GB of data used per month.

darkknight:

get to ya in a few min

darkknight:

Okay, so lets look at Company A. 50$ per month for unlimited minutes and texts. since this is per month, and we are finding the function per month. We can have the y-intercept in this function equal to 50. Now $1.5 per GB so this is our slope. SImilarly for Company B $35 per month for unlimited minutes and texts. since this is per month, and we are finding the function per month. We can have the y-intercept in this function equal to 35. Now $5 per BG so this is our slope for this function. Can you now write out the functions for both companies knowing the slopes and y-intercepts? Hint: use slope intercept form, where y = mx+b

zarkam21:

50=1.5x+35

zarkam21:

@darkknight

darkknight:

well we have 2 different functions. So the first one would have a slope of $1.5 and a y-inter of $50. The second function will have a slope of $5 and a y-inter of $35.

darkknight:

You have your slope and y-intercepts for BOTH functions so you can make 2 functions.

zarkam21:

50=1.5x 35=5x

zarkam21:

@Hero @Vocaloid

darkknight:

No @zarkam21 you have the slope. Do you know what the y-intercept is? y=mx+b?

zarkam21:

isn't it $50?

darkknight:

yes. Then why do you have that for y. Y intercept in y = mx+b is b

zarkam21:

y=1.5x+50 y=5x+35

darkknight:

Yes there you go. remember both are in dollars Keep up the good work

zarkam21:

thank you so much!

darkknight:

np I am the darkknight